A particle with velocity v0= -2i+4j in meters per second at t=0 undergoes a constant acceleration a of magnitude a=3ms^-2 at an angle =130° from the positive direction of the x axis. What is the particles velocity v at t=5s, in unit vector notation and as a magnitude and an angle?
1
Expert's answer
2019-12-16T11:05:19-0500

First of all we determine the unit vector with direction of acceleration. This may be done by expression "\hat n=\hat i\cdot cos\alpha+\hat j\cdot sin\alpha" , where "\alpha=130\degree". Vector of acceleration will be "\overrightarrow{a}=a\cdot \hat n" . The equation of evolution of velocity is "\overrightarrow{V}=\overrightarrow{V_0}+\overrightarrow{a} \cdot t" . In unit vector notation we can rewrite this equation in the form "\overrightarrow{V}=-2\hat i+4 \hat j+a\cdot t\cdot cos\alpha\cdot \hat i+a\cdot t\cdot sin\alpha\cdot \hat j" . After reduction of similar terms we get "\overrightarrow{V}=(-2+a\cdot t\cdot cos\alpha )\cdot\hat i+(4+a\cdot t\cdot sin\alpha) \hat j" . At "t=5s" the velocity will be

"\overrightarrow{V}[m/s]=(-2+3\cdot 5\cdot cos (130\degree) )\cdot\hat i+(4+3\cdot 5\cdot sin(130 \degree)) \hat j=" "-11.64\cdot \hat i+15.49\cdot \hat j"

It's magnitude is "V=\sqrt{(\overrightarrow{V}\cdot \overrightarrow{V})}=\sqrt{(-11.64)^2+(15.49)^2}=19.38\ m/s" and an angle from the positive direction of the x axis is "\beta=acos(-11.64/19.38)=126.91\degree" .

Answer: The partical velocity is "\overrightarrow{V}[m/s]=-11.64\cdot \hat i+15.49\cdot \hat j" the magnitude "19.38\ m/s" , and an angle from the positive direction of the x axis is "126.91\degree" .

.